OpenStudy (anonymous):

A bug walking on the circular rim of a flower pot completes one lap around the planter in 17.27 s. If the radius of the planter is 10.7 cm, how fast was it traveling?

OpenStudy (anonymous):

Hey, This can be solved like any average velocity question with a simple distance divided by time equation, To figure out the distance you need only calculate the circumference of the flower pot using the following equation: \[2\pi r = Circumference = Distance\] Where r is the radius. \[Distance = 2\times \pi \times0.107 metres = 0.67 metres\] Once you have your distance just divide by the time taken and this will give the speed, as shown below, \[Speed = Distance (m) \div Time (s) = 0.67m \div 17.27s = 0.04m/s \] Remember always to convert your units into SI units, such as centimetres into metres otherwise your calculations will be wrong. Also somethign physics/maths teachers love is when you derive an equation and only put your numbers in at the end, so you could also solve your problem like so, \[2\pi r \div Time = Speed\] Hope this helps